OpenStudy (anonymous):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 3/4?

TheSmartOne (thesmartone):

Hi! :)

TheSmartOne (thesmartone):

Slope-intercept form is y = mx + b where m is slope you are given a slope of 3/4 so plug that in first :)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

what do i do after that?

TheSmartOne (thesmartone):

so first plug it in and then I'll go over that :)

OpenStudy (anonymous):

i plugged it in.

TheSmartOne (thesmartone):

so what did you get? :)

OpenStudy (anonymous):

y=3/4x+b

OpenStudy (anonymous):

so what now?

TheSmartOne (thesmartone):

now we plug in the point (5,-4)

TheSmartOne (thesmartone):

that is the form of (x,y) so x = 5 y = -4 plug it in -4 = (3/4)(5) + b solve for b :)

TheSmartOne (thesmartone):

\(\sf\Large -4 = (\frac{3}{4})(5) + b\) Solve for b :)

OpenStudy (anonymous):

but the choices i have are: A)y=-3/4x-1/4 B)y=3/4x+1/4 C)y=3/4x-31/4 D)y=3/4x+31/4

TheSmartOne (thesmartone):

first solve it for b, and then we can get the answer